# Low-Cost Architecture and Efficient Pattern Synthesis for Polarimetric Phased Array Based on Polarization Coding Reconfigurable Elements

### v2

@@ IV Parameter Sensitivity Analysis
IV-A Polarization State
The desired polarization state affects the pattern performance by modifying the polarization vector’s decomposition results given by Equation (17), and further changing element numbers $n_{h}$, $n_{v}$, and compensation phase $\beta$.
Fig. 13 presents the proposed PCRPA and the CPPA pattern performance concerning different desired polarization states. The results are obtained concerning the synthesis of arbitrarily polarized patterns. The polarization state parameters are adjusted by varying either $\gamma$ (with $\eta$ fixed at $60^{\circ}$) or $\eta$ (with $\gamma$ fixed at $30^{\circ}$). First, compared to the curves shown in Fig. 13c and 13d, the PCRPA curves in Fig. 13a and 13b exhibit more pronounced fluctuations, indicating a stronger sensitivity to $\gamma$ than $\eta$. Then, considering the impact of $\gamma$, the XPL is consistently suppressed below -40 dB and exhibits little correlation with the polarization state. In contrast, the PSL, directivity, and power curves show a more regular trend that changes with $\gamma$. When $\gamma$ approaches $45^{\circ}$, the PSL and directivity become close to the CPPA values, but the power loss increases to a maximum of around 3 dB. This is because the number of H- and V-polarized elements $n_{h}$ and $n_{v}$ tend to be equal, resulting in similar array factors to those of the CPPA. Meanwhile, a significant portion of power is canceled during the vector composing process.
These can lead to the following conclusions. For a given array size and beam direction:
  • The XPL in PCRPA pattern synthesis is almost insensitive to the desired polarization state, while the PSL, directivity, and radiated power are sensitive to the auxiliary polarization angle $\gamma$.
  • The PCRPA exhibits lower PSL and directivity compared to the CPPA. Their performance difference gradually diminishes as $\gamma$ approaches $45^{\circ}$, until the characteristics become nearly comparable.
  • The PCRPA exhibits a certain level of power loss compared to the CPPA. The loss magnitude progressively increases as $\gamma$ approaches $45^{\circ}$, with the maximum loss reaching approximately 3 dB.
IV-B Beam Direction
Figure 14: Pattern performance of the proposed and conventional polarimetric phased arrays under different beam directions. In the legend, ’1’ represents the single-polarized beam, and ’2’ denotes the dual-polarized beam. (a) PSL curves. (b) XPL curves. (c) Directivity curves.
@@ IV-B Beam Direction
Figure 15: Pattern performance of the proposed and conventional polarimetric phased arrays under different array sizes. In the legend, ’1’ represents the single-polarized beam, and ’2’ denotes the dual-polarized beam. (a) PSL curves. (b) XPL curves. (c) Directivity curves.
The beam direction affects the pattern performance by determining the specific basis vector sets selected through Equation (15), as well as the element pattern data and weight vectors used in Equation (21).
Fig. 14 shows the performance of both single- and dual-polarized pattern synthesis of the PCRPA and the CPPA concerning different beam directions. The azimuth angle $\phi$ is fixed at $15^{\circ}$ while the elevation angle $\theta$ ranges from $30^{\circ}$ to $150^{\circ}$ with a step of $15^{\circ}$. The polarization state of the single-polarized pattern is fixed at $(30^{\circ},60^{\circ})$. Figs. 14a and 14c reveal the same degradation trends of PSL and directivity for both architectures concerning single- and dual-polarized cases with increasing scanning angles. Additionally, the performance gaps between the PCRPA and the CPPA remain stable throughout the scanning range. In Fig. 14b, the XPL of the single-polarized pattern appears insensitive to scanning angle variations. In contrast, the dual-beam XPL demonstrates an increasing trend as the scanning angle $\Delta\theta$ enlarges. When $\Delta\theta\leq 30^{\circ}$, the dual-beam XPL is lower than the single-beam by 10-20 dB; however, they converge to comparable values when $\Delta\theta>30^{\circ}$. This owes to the different XPL suppression mechanisms: the dual-beam employs the calibration matrix with continuously adjustable coefficients, while the single-beam relies on controlling the ratio of H- and V-polarized elements, where the element numbers are discrete. In addition, beam pointing error at large scanning angles induces a mismatch in the calibration matrix, deteriorating the XPL in dual-beam synthesis.
These can lead to the following conclusions. For a given array size and polarization state:
  • All pattern performance of the PCRPA, except the XPL of the single-polarized beam, exhibits sensitivity to beam scanning, with progressive degradation observed as the scanning angle increases.
  • The gaps in PSL and directivity between the PCRPA and CPPA remain almost stable throughout the scanning range.
  • The XPL of the PCRPA is always suppressed at no more than -40 dB, where the dual beam pattern synthesis can exhibit excessively low levels at -60 dB near the broadside.
IV-C Array Size
The array size affects the pattern performance by determining the specific number of H- and V-polarized elements in Equation (18), and the dimensionality of the optimization model given by Equation (27).
As shown in Fig. 15, all radiation pattern parameters exhibit evident improvements with the array size, except the XPL of dual-beam for the PCRPA. Notably, the XPL of dual-beam PCRPA consistently maintains -60 dB-level stability and demonstrates insensitivity to the array size. This is because its XP suppression performance is determined by the estimation accuracy of the polarization calibration matrix rather than the array size. Particularly, both the PSL of dual-beam synthesis and the XPL of single-beam synthesis show remarkable array size dependence in small-to-medium arrays ($\leq 24\times 24$): The performance enhancement in dual-beam synthesis stems from effective dimensionality expansion of the optimization model. In contrast, for the single-beam case, the improvement arises from the quantization characteristics of the polarization angle $\gamma$, given by the equation below:
@@ IV-C Array Size
Increasing the array size will reduce the quantization step of $\hat{\gamma}$, thereby improving polarization-state synthesis precision and enhancing XPL performance.
These can lead to the following conclusions. For a given polarization state and beam direction:
  • All PCRPA pattern performance, except for the dual-beam XPL, exhibits array-size dependence and shows progressive improvement with the array size.
  • The XPL performance of PCRPA dual-beam synthesis demonstrates negligible dependence on the array size.
  • In small-to-medium scale arrays, both the single-beam XPL and dual-beam PSL of the PCRPA demonstrate evident sensitivity to the array size, with rapid performance enhancement observed as the size expands. This scaling characteristic strongly suggests the proposed PCRPA architecture for large-scale PPA systems.
V Experimental Results and Discussion
@@ V-A Experimental Results
To verify the effectiveness of the proposed architecture of the PCRPA and synthesis methods, an experiment is carried out using an X-band $8\times 8$-unit dual-polarized microstrip patch array antenna, shown in Fig. 16. The operating frequency is 9.30 GHz. The measurement results include active element pattern data for all ports, as shown in Fig. 16b, which means that the whole array can be utilized as a fully digital system.
Figure 16: An X-band $8\times 8$-unit dual-polarized microstrip patch array antenna test in an anechoic chamber. (a) Front view. (b) Back view.
Fig. 17 shows the CoP and XP patterns under the conventional and proposed architecture for single polarized pattern synthesis, with the beam direction at $(115^{\circ},10^{\circ})$ and the polarization state at V polarization. Compared with the CPPA, the PCRPA pattern mitigates the XPL from -19.67 dB to -37.75 dB, improving approximately 18 dB. Although this value indicates a degraded performance compared to the figure of simulated cases at -40 dB given in Fig. 15b, it still agrees with the previous conclusion of XPL suppression. Besides, the values of PSL for both conventional and proposed architectures remain comparable at -13.40 dB and -13.99 dB, respectively.
@@
Figure 18: Measured patterns (elevation cut) under different architectures for dual polarized pattern synthesis. (a) Conventional PPA. (b) Proposed PCRPA.
Fig. 18 shows the 2-D power patterns measured under two architectures for dual-polarized pattern synthesis, with the beam direction the same as the previous single-beam case. For the CoP patterns, both architectures demonstrate comparable matching performance: CPPA achieves a 3dB beamwidth of $5.79^{\circ}$ for H- and V-polarized beams, while PCRPA attains a narrower $4.98^{\circ}$ in both polarizations. The corresponding matching errors measure 5.35dB (CPPA) and 9.26dB (PCRPA). Compared to the CPPA, the CoP patterns of PCRPAs show PSL increase at 1.76 dB and directivity loss at 2.65 dB due to fewer channels. In addition, the XPL is effectively suppressed by calibration methods from -19.88 dB to -25.20 dB. The experiment results agree with the conclusions drawn by simulations but also reveal performance deterioration due to practical factors.
@@ V-B Discussion
Analysis of both simulation and experimental data shows that the array size is a crucial factor for the performance of the proposed PCRPA architecture. Employing PCRPAs in larger arrays with a minimum size of $8\times 8$ is recommended. As the array size increases, the proposed PCRPAs exhibit lower sidelobe levels and better control over cross-polarization levels of the patterns. This can be attributed to several reasons:
@@
    For the dual polarized pattern synthesis, larger arrays can offer more DOFs in the thinned array synthesis problem, allowing for higher optimization limits on the PSL of the CoP patterns.
@@ VI Conclusion
Despite the advantages, simulation results show that the proposed technique exhibits sensitivity to array size in terms of PSL and XPL in the pattern synthesis. Especially for XPL, the mitigation effect is influenced by multiple factors, such as quantization errors and mutual coupling. Moreover, due to the reduction in the number of T/R channels, there is an inevitable loss in beam power and directivity. Future research will focus on these problems.

### v3

@@ IV-C Parameter Sensitivity Analysis
IV-C1 Polarization State
This experiment evaluates the impact of polarization parameters on synthesis performance. The polarization states for both the CPPA and PCRPA were synthesized by traversing either $\gamma$ or $\eta$ while keeping the other fixed at $\eta=60^{\circ}$ or $\gamma=30^{\circ}$, respectively, with subsequent pattern metrics collected. To further quantify the sensitivity to polarization parameters, a new variable $P_{t}$, defined as the radiated power of the CoP pattern at the beam direction, was introduced.
The results given by Fig. 11 reveal several key findings. First, the CPPA is largely insensitive to polarization variations, whereas the PCRPA exhibits significant sensitivity. Second, the PCRPA shows low sensitivity to the polarization angle $\eta$; although the XPL curve in Fig. 11e exhibits fluctuations, all values remain below -40 dB, and the irregular variations can reasonably be attributed to differences in active element patterns under mutual coupling. Crucially, the PCRPA’s performance metrics—PSL, Dir, and $P_{t}$—are highly sensitive to the polarization angle $\gamma$. As shown in Figs. 11b-11d, when $\gamma$ approaches $0^{\circ}$, $45^{\circ}$, or $90^{\circ}$, the patterns display higher PSL and Dir. However, a distinct $P_{t}$ loss of approximately 3 dB occurs near $\gamma=45^{\circ}$, while this loss diminishes to nearly zero as $\gamma$ approaches $0^{\circ}$ or $90^{\circ}$.
This behavior originates from the synthesis mechanism of the PCRPA: the ratio of H-to-V ports is determined by $\gamma$ given by Equation (19). When this ratio approaches 1, or when ports are entirely H or V, the PCRPA’s CoP pattern most closely resembles that of the CPPA, leading to similar PSL and Dir. However, the vector superposition of orthogonal radiation fields from a nearly equal number of ports results in energy cancellation during polarization synthesis, which causes the observed $P_{t}$ degradation. This warrants careful consideration in low SNR environments where power efficiency is critical for detection accuracy.
IV-C2 Array Size
@@
Figure 12: Comparison of pattern performance between the conventional and proposed arrays across different array sizes. The legend distinguishes the beam types: ’1’ for single-polarized and ’2’ for dual-polarized beams. (a) XPL. (b) PSL. (c) Directivity.
Figure 13: Comparison of pattern performance between the conventional and proposed arrays across different beam directions. The legend distinguishes the beam types: ’1’ for single-polarized and ’2’ for dual-polarized beams. (a) XPL. (b) PSL. (c) Directivity.
This experiment investigates the impact of array size on synthesis performance. With the beam direction and polarization fixed, arrays of sizes $8\times 8$, $16\times 16$, $24\times 24$, and $32\times 32$ were evaluated. The AEPs of these arrays served as the input for both CPPA and PCRPA in arbitrary-polarized and dual-polarized pattern synthesis, with subsequent pattern metrics calculated.
The results given by Fig. 12 reveals several key trends. First, as shown in Fig. 12a, the XPL of the PCRPA decreases rapidly as the array size increases from $8\times 8$ to $24\times 24$, plateauing after falling below -50 dB. Second, Fig. 12b indicates that the PSL for dual-beam synthesis in the PCRPA improves significantly when scaling from $8\times 8$ to $16\times 16$, then stabilizes, whereas its single-beam PSL declines gradually throughout. Furthermore, the directivity of both CPPA and PCRPA increases with array size, following similar growth trends.
These findings indicate that the XPL and dual-beam PSL of the PCRPA are sensitive to the array size. The sensitivity of the dual-beam PSL shares a similar principle with thinned array synthesis, primarily governed by the dimensionality of the optimization problem. The sensitivity of the XPL, however, is inherent to its suppression mechanism of the PCRPA. The polarization angle $\hat{\gamma}$ in the PCRPA is a discrete, quantized value when mutual coupling is neglected, which can be expressed by:
@@
A comparison of the $\hat{\gamma}$ quantization curves for the $8\times 8$ and $32\times 32$ arrays in Fig. 14 clearly shows that a larger array results in a finer quantization step. This enhances the accuracy of approximating the desired polarization state, thereby improving polarization purity and yielding a lower XPL.
Figure 14: The quantization curves of $\hat{\gamma}$ concerning different array size.
IV-C3 Beam Direction
This experiment evaluates the impact of beam direction. With the azimuth angle fixed at $\phi=30^{\circ}$, the elevation angle $\theta$ was swept from $30^{\circ}$ to $150^{\circ}$ to define the beam direction for both CPPA and PCRPA in arbitrary-polarized and dual-polarized pattern synthesis. The corresponding pattern metrics were recorded.
The results given by Fig. 13 reveal several key observations. First, the CPPA’s XPL, PSL, and directivity all gradually degrade with beam scanning. Second, as shown in Figs. 13b and 13c, the PSL and directivity of the PCRPA follow the same trend as the CPPA. Although the proposed algorithm maintains PSL at a lower level than the CPPA, a consistent directivity loss is observed—approximately 0.3 dB for single-beam and 2.5 dB for dual-beam synthesis. Finally, Fig. 13a demonstrates that the PCRPA achieves significant XPL suppression, with the improvement becoming more pronounced at larger scanning angles. The single-beam synthesis yields better XPL suppression compared to the dual-beam case; however, the XPL values themselves fluctuate without a clear correlation to the scanning angle.
The similarity in PSL and directivity trends between the PCRPA and CPPA is attributed to their shared dependence on the element pattern. The XPL suppression in single-beam synthesis occurs because all array DoFs are allocated to synthesizing a single polarization state, resulting in finer equivalent polarization port quantization. This principle aligns with the array size sensitivity analysis discussed previously.
V Experiment and Discussion
@@ V-A Experimental Results
To verify the effectiveness of the proposed architecture of the PCRPA and synthesis methods, an experiment is carried out using a standard commercial dual-polarized microstrip patch antenna, shown in Fig. 15. The antenna operates at a frequency of 9.3 GHz, with a bandwidth of 200 MHz. The array size of $8\times 8$, and the element spacing is 16.1 mm. A planar near-field measurement was conducted for the antenna. During the test, the polarized ports were connected via a switch matrix to a vector network analyzer through independent channels, as shown in Fig. 15b. For all these ports, the 3-D AEPs were acquired via probe scanning and near-field to far-field transformation as measured results, which were then used as the input in the simulation platform. By controlling the 0-1 coding and phase weighting corresponding to each data set, the physical functions of the RF switches and phase shifters given in Fig. 2 were emulated, thereby validating the principle of the proposed architecture and methodology.
Figure 15: A standard commercial X-band $8\times 8$-unit dual-polarized microstrip patch array antenna test in an anechoic chamber. (a) Front view. (b) Back view.
@@
Fig. 16 compares the CoP and XP patterns of the conventional and proposed architectures for single-polarized pattern synthesis, under the same beam direction $(120^{\circ},15^{\circ})$ and V-polarization setting as used in simulations. The measured results align well with the simulation trends: the PCRPA achieves a significant XPL suppression from -18.48 dB (CPPA) to -35.26 dB, an improvement of approximately 16.78 dB. While this value is slightly higher than the simulated -40 dB level in Fig. 12b, it still supports the conclusion that the proposed architecture effectively suppresses XPL. Furthermore, the PSL remains nearly identical between the two architectures (-13.40 dB for CPPA vs. -14.30 dB for PCRPA), and a limited directivity loss of 0.68 dB is observed in the PCRPA. These experimental findings—XPL suppression, maintained PSL, and a minor directivity trade-off—collectively validate the key performance trends predicted by the simulations.
Figure 17: Measured patterns (elevation cut) under different architectures for dual polarized pattern synthesis. (a) H-polarized beam. (b) V-polarized beam.
Fig. 17 presents the measured 2-D power patterns for dual-polarized synthesis under the conventional and proposed architectures, also with the beam direction set to $(120^{\circ},15^{\circ})$. The experimental outcomes confirm the expected trade-offs: the PCRPA effectively suppresses XPL, reducing it from -14.23/-18.48 dB to -25.62/-39.84 dB through element polarization port reconfiguration. However, as also indicated in simulations, this improvement is accompanied by a PSL increase of 3.25 dB and a more noticeable directivity loss of 3.66 dB, attributable to the reduction in active channels. It is observed that the H-polarized beam exhibits less XPL suppression than its V-polarized counterpart, a phenomenon we attribute to the heightened sensitivity of mutual coupling, fabrication and environmental variations in smaller arrays. It is anticipated that this performance would improve with larger array scales, as stated in the sensitivity analysis, where such practical factors exert a relatively smaller influence. Overall, these results align with simulated predictions while underscoring the impact of real-world constraints on achievable performance.
@@ V-B Discussion
Analysis of both simulation and experimental results shows that the array size is a crucial factor for the performance of the proposed PCRPA architecture. Employing PCRPAs in larger arrays is recommended. As the array size increases, the proposed PCRPAs exhibit lower sidelobe levels and better control over cross-polarization levels of the patterns. This can be attributed to several reasons:
@@
    For the dual polarized pattern synthesis, larger arrays can offer more DoFs in the thinned array synthesis problem, allowing for higher limits on the PSL of the CoP patterns.
@@ VI Conclusion
Despite the advantages, simulation results show that the proposed technique exhibits sensitivity to array size in terms of PSL and XPL in the pattern synthesis. Especially for XPL, the mitigation effect is influenced by multiple factors, such as quantization errors and mutual coupling. Moreover, due to the reduction in the number of T/R channels, there is an inevitable loss in beam power and directivity.
Looking forward, future research will address the identified limitations in PSL and XPL performance. Additionally, extending the proposed architecture to conformal arrays presents an important research direction. In such geometries, position-dependent polarization variations and complex mutual coupling effects due to surface curvature introduce significant challenges. Future work will explore modeling strategies and algorithmic adaptations to handle these problems.
[Uncaptioned image] Yiqing Wang was born in 1999. She received the B.S. degree in information engineering from the National University of Defense Technology, Changsha, China, in 2021, where she is currently pursuing the Ph.D. degree in information and communication engineering. Her research interests include antenna arrays and polarimetric phased arrays.
[Uncaptioned image] Jian Zhou was born in 1996. He received the B.S. degree in information engineering and the Ph.D. degree in information and communication engineering from the National University of Defense Technology (NUDT), Changsha, China, in 2018 and 2024, respectively. In 2025, he joined the College of Electronic Science, NUDT, where he is currently an Assistant Researcher. His research interests include antenna arrays and polarimetric phased array.
[Uncaptioned image] Chen Pang was born in 1986. He received the B.S. and Ph.D. degrees in information and communication engineering from the National University of Defense Technology (NUDT), Changsha, China, in 2009 and 2015, respectively. From November 2012 to November 2014, he was a visiting Ph.D. Student with the Delft University of Technology, Delft, The Netherlands. Currently, he is an Associate Professor with the College of Electronic Science and Technology, National University of Defense Technology. His research interests include weather radar signal processing, radar polarimetry, active array, and electromagnetic simulation.
[Uncaptioned image] Wenyang Man was born in 2000. He received the B.S. degree in Electronic Engineering from North China Electric Power University, Baoding, China, in 2023, and is currently pursuing the Ph.D. degree in Electronic Science and Technology at the National University of Defense Technology, Changsha, China. His research interests include shared-aperture polarimetric phased array radar.
[Uncaptioned image] Zixiang Xiong was born in 2000. He received the M.S. degree in information engineering from the National University of Defense Technology, Changsha, China, in 2025, where he is currently pursuing the Ph.D. degree in electronic information. His research interests include antenna arrays and polarimetric phased arrays.
[Uncaptioned image] Ke Meng was born in 1999. He received the B.S. degree in information engineering from the National University of Defense Technology, Changsha, China, in 2021, where he is currently pursuing the M.S. degree in electronic information. His research interests include antenna arrays and polarimetric phased arrays.
[Uncaptioned image] Zhanling Wang was born in 1988. He received the B.S. degree in electronic engineering from Shenyang University of Technology, Shenyang, China, in 2007, the M.S. degree in information and communication engineering from Air Force Engineering University, Xi’an, China, in 2011, and the Ph.D. degree from the National University of Defense Technology, Changsha, China, in 2021. He is currently a Lecturer. His research interests include array antennas design and polarimetric phased arrays.
[Uncaptioned image] Yongzhen Li was born in 1977. He received the B.S. and Ph.D. degrees in electronic engineering from the National University of Defense Technology, Changsha, China, in 1999 and 2004, respectively. Currently, he is a Professor with the National University of Defense Technology. His research interests include radar signal processing, radar polarimetry, and target recognition.